Please answer only the empty boxes (bottom table and the questions below it)

The total subtracting cost= $_______

The total inventory holding cost for January through August= $_____

The total​ cost, excluding normal time labor​ costs, for Plan E= $_____

Will thumps up if answers is correct!

The president of Hill Enterprises, Terri Hill, projects the firm's aggregate demand requirements over the next 8 months as follows: 1,400 1,700 May 2,300 January February 2,300 June March 1,700 1,700 July April 1,900 August 1.500 Her operations manager is considering a new plan, which begins in January with 200 units of inventory on hand. Stockout cost of lost sales any idle-time costs. Evaluate the following plans D and E. $70 per unit. Inventory holding cost is $25 per unit per month. Ignore the reqular production, another 20 % Plan Keep the current workforce stable at producing 1,600 units per month. In addition the normal production units can be produced in overtime at an additional cost of $55 per unit. A warehouse now constrains the maximum allowable inventory hand to 600 units or less. Note: Do not produce in overtime if production or inventory are adequate to cover demand
Note: Do not produce in overtime if production or inventory are adequate to cover demand. Plan D 0.T Production Ending Stockouts n onanpo (Units) (Units) Month Demand (Units) Inv ry 200 0 December 1 January 1,400 1,600 0 400 0 2 February 1.700 1,600 0 300 3 March 1,700 1,600 0 200 4 April 1,900 1,600 100 0 0 5 May 2,300 1,600 320 380 0 6 June 2,300 1,600 380 320 0 7 July 1,700 1,600 100 0 8 August 1,600 1,500 0 100 $ 46200. (Enter your response as a whole number.) The total overtime production cost The total inventory holding cost for January through August $ 25000. (Enter your response as a whole number.) The total stockout cost $53200. (Enter your response as a whole number.) a whole number.) The total cost, excluding normal time labor costs, for Plan D $ 124400. (Enter your response
Plan E: Keep the current workforce, which is producing 1,600 units per month, and subcontract month. The warehouse and overtime constraints from Plan D do not apply to this plan. meet the rest of the demand. Subcontract cost is $80 per unit. Subcontracting capacity is limited to 700 units per Plan E Production Subcontract Ending Inventory (Units) Month Demand (Units) 0 December 200 January 1,400 1,600 2 February 1,600 1,700 3 March 1,600 1.700 4 April 1.900 1,600 5 May 2.300 1,600 6 June 2.300 1,600 7 July 1,700 1,600 8 August 1.500 1.600
Answer #1

subcontract is required when production is less than demand

Maximum subcontract = 700

For period t,

Subcontractt = min(max(Demandt - Productiont - Ending Inventoryt-1, 0),700)

Ending Inventoryt = max(Ending Inventoryt-1 + Productiont - Demandt, 0)

Month Demand Production Subcontract Ending Inventory
0 200
1 January 1400 1600 0 400
2 February 1700 1600 0 300
3 March 1700 1600 0 200
4 April 1900 1600 100 0
5 May 2300 1600 700 0
6 June 2300 1600 700 0
7 July 1700 1600 100 0
8 August 1500 1600 0 100
Total 14500 12800 1600 1000

Total Subcontracting Cost = 1600*80 = $ 128000

Total Inventory Holding Cost = 1000*25 = $25000

Total Cost = 128000 + 25000 = $153000